My question is that if I have a triangle with the angles of x, 3x, and 6x and the longest side of the triangle is 7 inches what would the equation be to find x to
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

if I have a triangle with the angles of x, 3x, and 6x and the longest side of the triangle is 7 inches what would the equation be to find x to
------
The longest side is opposite the largest angle.
6x = 7
x = 7/6
---
So, 3x = 3(7/6) = 7/2
-------------
Cheers,
Stan H.
-------------

Answer by ikleyn(52748)

.
My question is that if I have a triangle with the angles of x, 3x, and 6x and the longest side of the triangle is 7 inches
what would the equation be to find x to
~~~~~~~~~~~~~~~~~~~~~~~~~

The sum of interior angles of ANY triangle is 180 degrees.


Therefore, your equation to find x is this:


x + 3x + 6x = 180  --->  10x = 180  --->  x = {{180/10}}} = 18.


Answer.  x = 18 degrees.


The info regarding 7 inches is not relevant to finding x.


This info is excessive and is not used in  the solution.